OpenStudy (anonymous):

A line has slope –5/3. Through which two points could this line pass? (1 point) (12, 13), (17, 10) (16, 15), (13, 10) (0, 7), (3, 10) (11, 13), (8, 18)

OpenStudy (anonymous):

slope=y2-y1/x2-x1 -5/3=18-13/8-11 option D

OpenStudy (anonymous):

What the... .-.

OpenStudy (anonymous):

yeah please stop tagging everyone i dont think anyone likes having like 23 notifications waiting and its just this stuff...

OpenStudy (hba):

I dont even know who the person is :/

OpenStudy (anonymous):

this post is not for question solving it is a gathering question of all solvers..........lol:)